The problem I am working on is :
3(x + 2) – 4x = 3

The way I was working it out is
3(x2)-4x=3
6x-4x=3
But I know this math isn't right.
What am I doing wrong?

Answer by Cromlix(4381)

3(x + 2) – 4x = 3
Multiply (x + 2) by 3
3x + 6 - 4x = 3
Collect like terms remembering to
change signs as you cross the = sign.
3x - 4x = 3 - 6
-x = -3
Multiply both sides by -1
x = 3
Proof:-
3(x + 2) – 4x = 3
3(3 + 2) - 4(3) = 3
3(5) - 12 = 3
15 - 12 = 3
Hope this helps.
:-)